OpenStudy (anonymous):

You are traveling on an airplane. The velocity of the plane with respect to the air is 110.0 m/s due east. The velocity of the air with respect to the ground is 40.0 m/s at an angle of 30° west of due north.1) What is the speed of the plane with respect to the ground?2) What is the heading of the plane with respect to the ground? (Let 0° represent due north, 90° represents due east).(East of due North).3) How far east will the plane travel in 1 hour?

OpenStudy (anonymous):

vpa=110 i vag=-40 cos30 i + 40 sin30 j 1) vpg=vpa+vag 2) the heading: theta=tan-1(y/x) angle with respect to east direction. 3) try find (1) and (2) first :D Is this helps you?

OpenStudy (anonymous):



OpenStudy (anonymous):

I found 1 and 2...how to i find 3?

OpenStudy (anonymous):

Velocity = Change in position / change in time

OpenStudy (anonymous):

i know that..thnks!

OpenStudy (anonymous):

We have to make sure here that you use vpg to find the distance. Not vpa :)

OpenStudy (anonymous):

I have to correct answer to 1, but I'm still having trouble finding 3. Can anyone help me?

OpenStudy (anonymous):

1. Vpa = 180m/s. @ 0 deg. Vag = 40m/s @ 120 deg,CCW. Vpg = Vpa + Vag, Vpg = (180 + 40cos120) + i40sin120, Vpg = 160 + i34.64, Vpg=sqrt((160)^2 + (34.64)^2)=163.7m/s. 2. tanA = Y / X = 34.64 / 160 = 0.2165, A = 12.2 deg,CCW. = 12.2deg. North of East. 3. 1 hr = 3600s. d = Vt = 163.7m/s * 3600s = 589,320m.

OpenStudy (anonymous):

Those are not the right answers